# Diffusion approximation of the stationary distribution of a two-level single server queue

### v4

@@ 5.1 Process limit
Because of the exponentially distributed assumption, for each fixed $r$, $L^{(r)}(\cdot)$ is a continuous-time Markov chain with transition rates $q^{(r)}_{i,j}$ for $i,j\in\mathbb{Z}_{+}$ given by
$q^{(r)}_{0,0}=-c^{(r)}_{1}\mu^{(r)},\quad q^{(r)}_{0,1}=\lambda^{(r)}_{1},$
$q^{(r)}_{i,i}=-(c^{(r)}_{1}\mu^{(r)}+\lambda^{(r)}_{1}),\quad q^{(r)}_{i,i-1}=c^{(r)}_{1}\mu^{(r)},\quad q^{(r)}_{i,i+1}=\lambda^{(r)}_{1},\quad 1\leq i\leq\ell^{(r)}_{0}-1,$
$q^{(r)}_{i,i}=-(c^{(r)}_{1}\mu^{(r)}+\lambda^{(r)}_{2}),\quad q^{(r)}_{i,i-1}=c^{(r)}_{1}\mu^{(r)},\quad q^{(r)}_{i,i+1}=\lambda^{(r)}_{2},\quad i=\ell^{(r)}_{0},$
$q^{(r)}_{i,i}=-(c^{(r)}_{2}\mu^{(r)}+\lambda^{(r)}_{2}),\quad q^{(r)}_{i,i-1}=c^{(r)}_{2}\mu^{(r)},\quad q^{(r)}_{i,i+1}=\lambda^{(r)}_{2},\quad i\geq\ell^{(r)}_{0}+1,$
where all the other $q_{i,j}$’s vanish. Thus, $L^{(r)}(\cdot)$ is a birth-and-death process. By the stability condition (3.e), it has the stationary distribution. Denote it by $\pi^{(r)}_{n}=\mathbb{P}(L^{(r)}=n)$ for $n\geq 0$, then it is easy to see that
$\pi^{(r)}_{n}=\left\{\begin{array}[]{ll}\pi^{(r)}_{0}(\rho^{(r)}_{1})^{n},&0\leq n\leq\ell^{(r)}_{0},\\ \pi^{(r)}_{0}(\rho^{(r)}_{1})^{\ell^{(r)}_{0}}(\rho^{(r)}_{2})^{n-\ell^{(r)}_{0}},&n>\ell^{(r)}_{0},\end{array}\right.$ (A.3)
where
$\pi^{(r)}_{0}=\left\{\begin{array}[]{ll}\displaystyle\frac{(1-\rho^{(r)}_{1})(1-\rho^{(r)}_{2})}{1-\rho^{(r)}_{2}+(\rho^{(r)}_{2}-\rho^{(r)}_{1})(\rho^{(r)}_{1})^{\ell^{(r)}_{0}}},&\rho^{(r)}_{1}\not=1,\\ \displaystyle\frac{1-\rho^{(r)}_{2}}{(\ell^{(r)}_{0}+1)(1-\rho^{(r)}_{2})+\rho^{(r)}_{2}},&\rho^{(r)}_{1}=1.\end{array}\right.$ (A.6)
Hence, we can directly compute $\widetilde{\nu}_{i}(\mathbb{R}_{+})$ for $i=1,2$ through $\pi^{(r)}_{0}$ of (A.6), which yields
$\nu^{(r)}_{1}(\mathbb{R}_{+})=\sum_{n=0}^{\ell^{(r)}_{0}}\pi^{(r)}_{n}=\left\{\begin{array}[]{ll}\displaystyle\frac{1-(\rho^{(r)}_{1})^{\ell^{(r)}_{0}+1}}{1-\rho^{(r)}_{1}}\pi^{(r)}_{0},&\rho^{(r)}_{1}\not=1,\\ \displaystyle(\ell^{(r)}_{0}+1)\pi^{(r)}_{0},&\rho^{(r)}_{1}=1,\end{array}\right.$ (A.9)
$\nu^{(r)}_{2}(\mathbb{R}_{+})=\sum_{n=\ell^{(r)}_{0}+1}^{\infty}\pi^{(r)}_{n}=\frac{\rho^{(r)}_{2}}{1-\rho^{(r)}_{2}}(\rho^{(r)}_{1})^{\ell^{(r)}_{0}}\pi^{(r)}_{0}.$ (A.10)
We now consider the asymptotic behaviors of $\pi^{(r)}_{0}$ and $\widetilde{\nu}_{i}(\mathbb{R}_{+})$ for $i=1,2$. Form the definition of $\rho^{(r)}_{i}$ and the assumptions (3.b) and (3.c),
$\rho^{(r)}_{i}=\frac{\lambda^{(r)}_{i}}{c^{(r)}_{i}\mu^{(r)}}=1-\frac{r\mu b_{i}+o(r)}{c^{(r)}_{i}\mu^{(r)}}=1-\frac{b_{i}}{c_{i}}r+o(r),\qquad r\downarrow 0.$
We first consider the case that $b_{1}\not=0$. Since $\beta_{i}=b_{i}/c_{i}$ by the exponential assumption,
$(\rho^{(r)}_{1})^{\ell^{(r)}_{0}}=\big{(}1-\beta_{1}r+o(r)\big{)}^{r^{-1}\ell_{0}+o(1)}=e^{-\beta_{1}\ell_{0}}+o(1)),$ (A.11)
and therefore (A.6) yields
$\pi^{(r)}_{0}$ $=\frac{\beta_{1}\beta_{2}r^{2}+o(r^{2})}{\beta_{2}r+o(r)+(e^{-\beta_{1}\ell_{0}}+o(1))((\beta_{1}-\beta_{2})r+o(r))}$
$=\frac{\beta_{1}\beta_{2}}{\beta_{2}+(\beta_{1}-\beta_{2})e^{-\beta_{1}\ell_{0}}}(r+o(r))=\frac{b_{1}}{c_{1}}\frac{c_{1}b_{2}e^{\beta_{1}\ell_{0}}}{c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)+c_{2}b_{1}}r+o(r),$ (A.12)
which is compatible with (4.59). It follows from (A.9) and (A.2) that
$\lim_{r\downarrow 0}\nu^{(r)}_{1}(\mathbb{R}_{+})$ $=\lim_{r\downarrow 0}\frac{1-e^{-\beta_{1}\ell_{0}}+o(1)}{\beta_{1}r+o(r)}\frac{\beta_{1}\beta_{2}}{\beta_{2}+(\beta_{1}-\beta_{2})e^{-\beta_{1}\ell_{0}}}(r+o(r))$
$=\frac{\beta_{2}(1-e^{-\beta_{1}\ell_{0}})}{\beta_{2}+(\beta_{1}-\beta_{2})e^{-\beta_{1}\ell_{0}}}=\frac{c_{1}b_{2}(e^{(b_{1}/c_{1})\ell_{0}}-1)}{c_{1}b_{2}(e^{(b_{1}/c_{1})\ell_{0}}-1)+c_{2}b_{1}},$ (A.13)
[…]

### v5

@@ 5.1 Process limit
We prove Proposition 3.1 only for $b_{1}\not=0$ because the case for $b_{1}=0$ can be similarly proved. We first consider this 2-level $M/M/1$ queue without index $r$ for notational convenience. Because of the exponential assumptions, we can describe this model by the continuous time Markov chain with state space
$S=\{(m,i);m=1,2,i=0,1,\ldots,\ell_{0}\}\cup\{(3,i);i\geq\ell_{0}+1\},$
and its transition diagram is depicted in Figure A.1.
Figure A.1: Transition diagram
Denote this Markov chain by $X_{\exp}(\cdot)\equiv\{(J(t),L(t));t\geq 0\}$. Note that arrivals may find the queue length less than $\ell_{0}$ when their last arrivals found the queue length not less than $\ell_{0}$ (see the arrival at time $t^{(r)}_{e}(8)$ in Figure 3.1). Because of this, the arrival rate at any time may depend on the queue length at the last arrival instant, and therefore $L(\cdot)$ cannot be a continuous-time Markov chain.
We partition the state space $S$ of $X_{\exp}(\cdot)$ to the following three disjoint subsets.
$S_{m}=\{(m,i)\in S;i\in[0,\ell_{0}]\},\quad m=1,2,\qquad S_{3}=\{(3,i):i\geq\ell_{0}+1\}.$
By the assumptions of Proposition 3.1, $X_{M}(\cdot)$ has a stationary distribution. Let $(J,L)$ be a random vector subject to this stationary distribution, and define $p_{m,i}$ as
$p_{m,i}=\mathbb{P}[J=m,L=i],\qquad m=1,2,3,(m,i)\in S_{m}.$
Obviously, the stationary distribution, $\{p_{m,i};(m.i)\in S\}$, is uniquely determined by the following stationary equations. For state in $S_{1}$,
$\lambda_{1}p_{1,0}=c_{1}\mu p_{1,1},$ (A.1)
$(\lambda_{1}+c_{1}\mu)p_{1,i}=\lambda_{1}p_{1,i-1}+c_{1}\mu p_{1,i+1}+\lambda_{2}p_{2,i-1},\quad i\in[1,\ell_{0}-1],$ (A.2)
$(\lambda_{1}+c_{1}\mu)p_{1,\ell_{0}}=\lambda_{1}p_{1,\ell_{0}-1}+\lambda_{2}p_{2,\ell_{0}-1},$ (A.3)
and, for states in $S_{2}$,
$\lambda_{2}p_{2,0}=c_{1}\mu p_{2,1},$ (A.4)
$(\lambda_{2}+c_{1}\mu)p_{2,i}=c_{1}\mu p_{2,i+1},\qquad i\in[1,\ell_{0}-1],$ (A.5)
$(\lambda_{2}+c_{1}\mu)p_{2,\ell_{0}}=c_{2}\mu p_{3,\ell_{0}+1},$ (A.6)
For state in $S_{3}$, it follows from the equation for $\ell_{0}+1$ and the local balance that
$(\lambda_{2}+c_{2}\mu)p_{3,\ell_{0}+1}=\lambda_{1}p_{1,\ell_{0}}+\lambda_{2}p_{2,\ell_{0}}+c_{2}\mu p_{\ell_{0}+2},$ (A.7)
$\lambda_{2}p_{3,i}=c_{2}\mu p_{3,i+1},\qquad i\geq\ell_{0}+1.$ (A.8)
In what follows, we use the following notations.
$\rho_{ij}=\frac{\lambda_{i}}{c_{j}\mu},\qquad\gamma_{ij}=\frac{c_{i}\mu}{c_{1}\mu+\lambda_{j}},\qquad i,j=1,2.$
We first note that, from (A.8),
$p_{3,i}=\rho_{22}^{i-(\ell_{0}+1)}p_{3,\ell_{0}+1},\qquad i\geq\ell_{0}+1.$ (A.9)
Thus, $p_{3,i}$ for $i\geq\ell_{0}+1$ is simply determined by $p_{3,\ell_{0}+1}$.
We next consider $p_{2,i}$ for $i\in[0,\ell_{0}-1]$. From (A.4) and (A.5), we have
$p_{2,0}=\frac{c_{1}\mu}{\lambda_{2}}p_{2,1}=\rho_{21}^{-1}p_{2,1},$ (A.10)
$p_{2,i}=\gamma_{12}p_{2,i+1}=\gamma_{12}^{\ell_{0}-i}p_{2,\ell_{0}},\quad i\in[1,\ell_{0}-1].$ (A.11)
Thus, $p_{2,i}$ for $i\in[0,\ell_{0}-1]$ is determined by $p_{2,\ell_{0}}$. Furthermore, (A.11) can be written as
$p_{2,i}=\gamma_{12}^{-1}p_{2,i-1}=\gamma_{12}^{-i+1}p_{2,1}=\gamma_{12}^{-i+1}\rho_{21}p_{2,0},\quad i\in[1,\ell_{0}-1].$ (A.12)
Let us consider relations among the stationary probabilities, $p_{1,0},p_{1,\ell_{0}}$, $p_{2,\ell_{0}}$ and $p_{3,\ell_{0}+1}$ at the boundaries states of $S_{m}$ for $m=1,2,3$. We call them boundary probabilities. There is one more boundary probability, $p_{2,0}$, but this probability is ignorable compared with $p_{2,\ell_{0}}$ for large $\ell_{0}$ as $\rho_{21}p_{2,0}=\gamma_{12}^{\ell_{0}-1}p_{2,\ell_{0}-1}=\gamma_{12}^{\ell_{0}}p_{2,\ell_{0}}$ by (A.12) and (A.11), so we do not consider it. We first note that, from (A.6),
$p_{2,\ell_{0}}=\gamma_{22}p_{3,\ell_{0}+1}.$ (A.13)
Since $c_{2}\mu p_{\ell_{0}+2}=\lambda_{2}p_{3,\ell_{0}+1}$ by (A.8) for $i=\ell_{0}+1$, (A.7) can be written as
$p_{3,\ell_{0}+1}=\rho_{12}p_{1,\ell_{0}}+\rho_{22}p_{2,\ell_{0}},$ (A.14)
and, substituting $p_{3,\ell_{0}+1}=\gamma_{22}^{-1}p_{2,\ell_{0}}$ of (A.13) to this equation, we have
$(\gamma_{22}^{-1}-\rho_{22})p_{2,\ell_{0}}=\rho_{12}p_{1,\ell_{0}}.$
Since $\gamma_{22}^{-1}-\rho_{22}=\frac{c_{1}\mu+\lambda_{2}}{c_{2}\mu}-\frac{\lambda_{2}}{c_{2}\mu}=\frac{c_{1}}{c_{2}}=\rho_{12}\rho_{11}^{-1}$, this equation yields
$p_{2,\ell_{0}}=\rho_{11}p_{1,\ell_{0}}.$ (A.15)
Thus, the boundary probabilities $p_{1,\ell_{0}}$, $p_{2,\ell_{0}}$ and $p_{3,\ell_{0}+1}$ are explicitly related by (A.13) and (A.15). We also like to see such a relation between $p_{1,0}$ and $p_{1,\ell_{0}}$, but this is not easy.
To resolve this problem, we consider $p_{1,i}$ for $i\in[0,\ell_{0}-1]$. From (A.2),
$\lambda_{1}p_{1,i}-c_{1}\mu p_{1,i+1}=\lambda_{1}p_{1,i-1}-c_{1}\mu p_{1,i}+\lambda_{2}p_{2,i-1},\qquad i\in[1,\ell_{0}-1],$
then this equation can be written as
$\rho_{11}p_{1,i}-p_{1,i+1}=\rho_{11}p_{1,i-1}-p_{1,i}+\rho_{21}p_{2,i-1},\qquad i\in[1,\ell_{0}-1].$
Summing up both sides of this equation for $i\in[1,j]$ for $j\in[1,\ell_{0}-1]$, we have
$\rho_{11}p_{1,j}-p_{1,j+1}=\rho_{21}\sum_{i=1}^{j}p_{2,i-1},\qquad j\in[0,\ell_{0}-1],$ (A.16)
because $\rho_{11}p_{1,0}-p_{1,1}=0$ by (A.1). This equation is valid also for $j=0$ by the convention that the empty sum vanishes, which we always take.
Hence, applying (A.10) and (A.11) to (A.16), we have
$\rho_{11}p_{1,j}-p_{1,j+1}$ $=\rho_{21}\left(\rho_{21}^{-1}p_{2,1}+\sum_{i=1}^{j-1}\gamma_{12}^{\ell_{0}-i}p_{2,\ell_{0}}\right)$
$=\left(1+\rho_{21}\sum_{i=1}^{j-1}\gamma_{12}^{-(i-1)}\right)\gamma_{12}^{\ell_{0}-1}p_{2,\ell_{0}}$
$=\left(1+\rho_{21}\frac{(\gamma_{12}^{-(j-1)}-1)\gamma_{12}}{1-\gamma_{12}}\right)\gamma_{12}^{\ell_{0}-1}p_{2,\ell_{0}},\qquad j\in[1,\ell_{0}-1].$
Multiplying both sides by $\rho_{11}^{-(j+1)}$,
$\rho_{11}^{-j}p_{1,j}-\rho_{11}^{-(j+1)}p_{1,j+1}$
$\quad=\rho_{11}^{-(j+1)}\left(1+\rho_{21}\frac{(\gamma_{12}^{-(j-1)}-1)\gamma_{12}}{1-\gamma_{12}}1(j\geq 1)\right)\gamma_{12}^{\ell_{0}-1}p_{2,\ell_{0}},\qquad j\in[0,\ell_{0}-1].$
Summing up both sides of this equation from $j=0$ to $j=k-1\leq\ell_{0}-1$, we have
$p_{1,0}-\rho_{11}^{-k}p_{1,k}=\sum_{j=0}^{k-1}\rho_{11}^{-(j+1)}\gamma_{12}^{\ell_{0}-1}p_{2,\ell_{0}}+\sum_{j=1}^{k-1}\frac{(\rho_{11}\gamma_{12})^{-(j-1)}-\rho_{11}^{-(j-1)}}{1-\gamma_{12}}\rho_{11}^{-2}\rho_{21}\gamma_{12}^{\ell_{0}}p_{2,\ell_{0}}$
$\quad=\frac{\rho_{11}^{-k}-1}{1-\rho_{11}}\gamma_{12}^{\ell_{0}-1}p_{2,\ell_{0}}+\left(\frac{(\rho_{11}\gamma_{12})^{-(k-1)}-1}{(\rho_{11}\gamma_{12})^{-1}-1}-\frac{\rho_{11}^{-(k-1)}-1}{\rho_{11}^{-1}-1}\right)\frac{\rho_{11}^{-2}\rho_{21}\gamma_{12}^{\ell_{0}}p_{2,\ell_{0}}}{1-\gamma_{12}}$
$\quad=\frac{\rho_{11}^{-k}-1}{1-\rho_{11}}\gamma_{12}^{\ell_{0}-1}p_{2,\ell_{0}}+\left(\frac{(\rho_{11}\gamma_{12})^{-(k-1)}-1}{(\rho_{11}\gamma_{12})^{-1}-1}\rho_{11}^{-2}-\frac{\rho_{11}^{-k}-\rho_{11}^{-1}}{1-\rho_{11}}\right)\frac{\rho_{21}\gamma_{12}^{\ell_{0}}p_{2,\ell_{0}}}{1-\gamma_{12}},\quad k\in[1,\ell_{0}].$
Hence, multiplying both sides of this equation by $\rho_{11}^{k}$, we have
$p_{1,k}$ $=\rho_{11}^{k}p_{1,0}-\frac{1-\rho_{11}^{k}}{1-\rho_{11}}\gamma_{12}^{\ell_{0}-1}p_{2,\ell_{0}}+\frac{1-\rho_{11}^{k-1}}{1-\rho_{11}}\frac{\rho_{21}}{1-\gamma_{12}}\gamma_{12}^{\ell_{0}}p_{2,\ell_{0}}$
$\quad-\frac{\gamma_{12}^{-(k-1)}-\rho_{11}^{k-1}}{(\rho_{11}\gamma_{12})^{-1}-1}\frac{\rho_{11}^{-1}\rho_{21}}{1-\gamma_{12}}\gamma_{12}^{\ell_{0}}p_{2,\ell_{0}},\qquad k\in[1,\ell_{0}].$ (A.17)
In particular, for $k=\ell_{0}$, substituting $p_{2,\ell_{0}}=\rho_{11}p_{1,\ell_{0}}$ of (A.15), we have
$\rho_{11}^{\ell_{0}}\frac{p_{1,0}}{p_{1,\ell_{0}}}$ $=1+\frac{1-\rho_{11}^{\ell_{0}}}{1-\rho_{11}}\gamma_{12}^{\ell_{0}-1}\rho_{11}-\frac{1-\rho_{11}^{\ell_{0}-1}}{1-\rho_{11}}\frac{\rho_{11}\rho_{21}}{1-\gamma_{12}}\gamma_{12}^{\ell_{0}}$
$\quad+\frac{\rho_{11}\gamma_{12}-\rho_{11}^{\ell_{0}}\gamma_{12}^{\ell_{0}}}{(\rho_{11}\gamma_{12})^{-1}-1}\frac{\rho_{11}^{-1}\rho_{21}}{1-\gamma_{12}}.$ (A.18)
Thus, $p_{1,0}$ is explicitly related to $p_{1,\ell_{0}}$.
We now apply these computations to the $r$-th 2-level $M/M/1$ queue, and compute the limit of its scaled stationary distribution through moment generating functions. For the $r$-th system, we put the superscript (r) as we have used for the $r$-th 2-level $GI/G/1$ queue. For example, $p^{(r)}_{m,i}$ is the stationary probability of $\{(J,L)=(m,i)\}$ for $(m,i)\in S$.
Define the moment generating functions $g^{(r)}_{m}(\theta)$ for $m=1,2,3$ as
$g^{(r)}_{m}(\theta)=\sum_{k=0}^{\ell^{(r)}_{0}}e^{r\theta k}p^{(r)}_{m,k},\quad\theta\in\mathbb{R},m=1,2,\qquad g^{(r)}_{3}(\theta)=\sum_{k=\ell^{(r)}_{0}+1}^{\infty}e^{r\theta k}p^{(r)}_{3,k},\quad\theta\leq 0,$
and define $g^{(r)}(\theta)=\sum_{m=1}^{3}g^{(r)}_{m}(\theta)$. Obviously, $g^{(r)}(\theta)$ is the moment generating function of $rL^{(r)}$. Thus, our task is to compute the limit of $g^{(r)}(\theta)$ as $r\downarrow 0$. To this end, we compute $g^{(r)}_{m}(\theta)$ for $m=1,2,3$, using the normalization condition that $g^{(r)}(0)=1$. However, it should be noted that this condition is not sufficient to prove the weak convergence of the distribution of $rL^{(r)}$, for which it is required that $\lim_{\theta\uparrow 0}\lim_{r\downarrow 0}g^{(r)}(\theta)=1$. This weak convergence will be verified through explicitly computing $g^{(r)}_{m}(\theta)$ for $m=1,2,3$.
We first compute $g^{(r)}_{m}(\theta)$ for $m=2,3$ because they are easier. It is immediate from (A.11), (A.10) and (A.9) that
$g_{2}^{(r)}(\theta)$ $=p^{(r)}_{2,0}+e^{\ell^{(r)}_{0}r\theta}\sum_{i=1}^{\ell^{(r)}_{0}}e^{-(\ell^{(r)}_{0}-i)r\theta}(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}-i}p^{(r)}_{2,\ell^{(r)}_{0}}$
$=\left((\rho^{(r)}_{21})^{-1}(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}-1}+e^{r\theta}\frac{e^{\ell^{(r)}_{0}r\theta}-(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}}{e^{r\theta}-\gamma^{(r)}_{12}}\right)p^{(r)}_{2,\ell^{(r)}_{0}},$ (A.19)
$g^{(r)}_{3}(\theta)$ $=\frac{1}{1-e^{r\theta}\rho^{(r)}_{2}}e^{(\ell^{(r)}_{0}+1)r\theta}p^{(r)}_{3,\ell^{(r)}_{0}+1}.$ (A.20)
On the other hand, computing $g^{(r)}_{1}(\theta)$ is a bit complicated. To ease it, we introduce the following functions.
$f^{(r)}_{1}(\theta)=\sum_{k=0}^{\ell^{(r)}_{0}-1}e^{r\theta k},\qquad f^{(r)}_{\rho_{1}}(\theta)=\sum_{k=0}^{\ell^{(r)}_{0}-1}(e^{r\theta}\rho^{(r)}_{1})^{k},\qquad f^{(r)}_{\gamma^{-1}_{12}}(\theta)=\sum_{k=0}^{\ell^{(r)}_{0}-1}(e^{r\theta}(\gamma^{(r)}_{12})^{-1})^{k},$
then
$f^{(r)}_{1}(\theta)=\frac{1-(e^{r\theta})^{\ell^{(r)}_{0}}}{1-e^{r\theta}}1(\theta\not=0)+\ell^{(r)}_{0}1(\theta=0),\qquad f^{(r)}_{\rho_{1}}(\theta)=\frac{1-(e^{r\theta}\rho^{(r)}_{1})^{\ell^{(r)}_{0}}}{1-e^{r\theta}\rho^{(r)}_{1}},$
$f^{(r)}_{\gamma^{-1}_{12}}(\theta)=\frac{e^{r\theta\ell^{(r)}_{0}}-(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}}{e^{r\theta}-\gamma^{(r)}_{12}}(\gamma^{(r)}_{12})^{-\ell^{(r)}_{0}+1}.$
Using these function, we compute $g^{(r)}_{1}(\theta)$ from (A.2) as given below.
$g^{(r)}_{1}(\theta)$ $=\sum_{k=0}^{\ell^{(r)}_{0}}(e^{r\theta}\rho^{(r)}_{1})^{k}p^{(r)}_{1,0}-\sum_{k=1}^{\ell^{(r)}_{0}}\frac{e^{r\theta k}-(e^{r\theta}\rho^{(r)}_{1})^{k}}{1-\rho^{(r)}_{1}}(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}-1}p^{(r)}_{2,\ell^{(r)}_{0}}$
$\quad+\sum_{k=1}^{\ell^{(r)}_{0}}\frac{e^{r\theta(k-1)}-(e^{r\theta}\rho^{(r)}_{1})^{k-1}}{1-\rho^{(r)}_{1}}\frac{e^{r\theta}\rho^{(r)}_{21}}{1-\gamma^{(r)}_{12}}(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}p^{(r)}_{2,\ell^{(r)}_{0}}$
$\quad-\sum_{k=1}^{\ell^{(r)}_{0}}\frac{(e^{r\theta}(\gamma^{(r)}_{12})^{-1})^{k-1}-(e^{r\theta}\rho^{(r)}_{1})^{k-1}}{(\rho^{(r)}_{1}\gamma^{(r)}_{12})^{-1}-1}\frac{e^{r\theta}(\rho^{(r)}_{1})^{-1}\rho^{(r)}_{21}}{1-\gamma^{(r)}_{12}}(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}p^{(r)}_{2,\ell_{0}}$
$=\left(f^{(r)}_{\rho_{1}}(\theta)+(e^{r\theta}\rho^{(r)}_{1})^{\ell^{(r)}_{0}}\right)p^{(r)}_{1,0}-\left(f^{(r)}_{1}(\theta)-\rho^{(r)}_{1}f^{(r)}_{\rho_{1}}(\theta)\right)\frac{e^{r\theta}(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}-1}}{1-\rho^{(r)}_{1}}p^{(r)}_{2,\ell^{(r)}_{0}}$
$\quad+\left(f^{(r)}_{1}(\theta)-f^{(r)}_{\rho_{1}}(\theta)\right)\frac{e^{r\theta}\rho^{(r)}_{21}}{(1-\rho^{(r)}_{1})(1-\gamma^{(r)}_{12})}(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}p^{(r)}_{2,\ell^{(r)}_{0}}$
$\quad-\left(f^{(r)}_{\gamma^{-1}_{12}}(\theta)-f^{(r)}_{\rho_{1}}(\theta)\right)\frac{e^{r\theta}(\rho^{(r)}_{1})^{-1}\rho^{(r)}_{21}}{((\rho^{(r)}_{1}\gamma^{(r)}_{12})^{-1}-1)(1-\gamma^{(r)}_{12})}(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}p^{(r)}_{2,\ell^{(r)}_{0}}$ (A.21)
We now compute the limits of $g^{(r)}_{m}(0)$ for $m=1,2,3$. For this, we note that the following asymptotic properties are obtained from the assumptions (3.b)–(3.e),
$(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}=O(\gamma_{12}^{\ell_{0}/r}),\qquad\mbox{as }r\downarrow 0.$ (A.22)
Hence, $(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}$ decays exponentially fast as the function of $1/r$. On the other hand,
$\lim_{r\downarrow 0}(1-\rho^{(r)}_{i})/r=\beta_{i},\qquad i=1,2,$ (A.23)
$\lim_{r\downarrow 0}(\rho^{(r)}_{1})^{\ell^{(r)}_{0}}=\lim_{r\downarrow 0}\left(1-\beta_{1}r+o(r)\right)^{\ell_{0}/r+o(r)}=e^{-\beta_{1}\ell_{0}},$ (A.24)
where recall that $\beta_{i}=b_{i}/c_{i}$. Another important asymptotic properties are the ratios of the boundary probabilities. They are presented in the next lemma.
Lemma A.1.
For the sequence of the 2-level $M/M/1$ queue satisfying (3.b)–(3.e),
$\lim_{r\downarrow 0}\frac{p^{(r)}_{2,\ell_{0}}}{p^{(r)}_{1,\ell_{0}}}=1,\qquad\lim_{r\downarrow 0}\frac{p^{(r)}_{3,\ell_{0}+1}}{p^{(r)}_{1,\ell_{0}}}=\frac{\lambda_{1}+\lambda_{2}}{\lambda_{2}},\qquad\lim_{r\downarrow 0}\frac{p^{(r)}_{1,0}}{p^{(r)}_{1,\ell_{0}}}=\frac{\lambda_{1}+\lambda_{2}}{\lambda_{2}}e^{\beta_{1}\ell_{0}}.$ (A.25)
Proof.
The first two limits in (A.25) aer immediate from (A.15) and (A.13) because
$\lim_{r\downarrow 0}\gamma^{(r)}_{22}=\lim_{r\downarrow 0}\frac{c^{(r)}_{2}\mu^{(r)}}{c^{(r)}_{1}\mu+\lambda^{(r)}_{2}}=\frac{c_{2}\mu}{c_{1}\mu+\lambda_{2}}=\frac{\lambda_{2}}{\lambda_{1}+\lambda_{2}}.$
For the third limit, we have, from (A.2) for the $r$-th system,
$\lim_{r\downarrow 0}\frac{p^{(r)}_{1,0}}{p^{(r)}_{1,\ell_{0}}}$ $=\lim_{r\downarrow 0}(\rho^{(r)}_{1})^{-\ell^{(r)}_{0}}\left(1+\frac{\rho^{(r)}_{1}\gamma^{(r)}_{12}}{(\rho^{(r)}_{1}\gamma^{(r)}_{12})^{-1}-1}\frac{(\rho^{(r)}_{1})^{-1}\rho^{(r)}_{21}}{1-\gamma^{(r)}_{12}}\right).$
Hence, applying (A.24), we have the third limit of (A.25). ∎
We now tentatively assume that there exists finite $\widetilde{p}_{1,\ell_{0}}$ such that
$\widetilde{p}_{1,\ell_{0}}=\lim_{r\downarrow 0}\frac{p^{(r)}_{1,\ell^{(r)}_{0}}}{r}.$ (A.26)
The existence of this $\widetilde{p}_{1,\ell_{0}}$ will be verified at the final step (see (A.31)). Applying these asymptotic properties, it follows from (A.2) and (A.20) with help of (A.26) that
$\lim_{r\downarrow 0}g_{2}^{(r)}(0)$ $=\lim_{r\downarrow 0}\left((\rho^{(r)}_{21})^{-1}(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}-1}+\frac{1-(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}}{1-\gamma^{(r)}_{12}}\right)p^{(r)}_{2,\ell^{(r)}_{0}}$
$=\frac{c_{1}\mu+\lambda_{2}}{\lambda_{2}}\lim_{r\downarrow 0}p^{(r)}_{2,\ell^{(r)}_{0}}=\frac{\lambda_{1}+\lambda_{2}}{\lambda_{2}}\lim_{r\downarrow 0}p^{(r)}_{1,\ell^{(r)}_{0}}=0,$ (A.27)
$\lim_{r\downarrow 0}g^{(r)}_{3}(0)$ $=\lim_{r\downarrow 0}\frac{1}{1-\rho^{(r)}_{2}}p^{(r)}_{3,\ell^{(r)}_{0}+1}=\frac{c_{2}}{b_{2}}\lim_{r\downarrow 0}\frac{1}{r}p^{(r)}_{3,\ell^{(r)}_{0}+1}$
$=\beta_{2}^{-1}\gamma_{22}^{-1}\lim_{r\downarrow 0}\frac{1}{r}p^{(r)}_{2,\ell^{(r)}_{0}}=\frac{\lambda_{1}+\lambda_{2}}{\lambda_{2}}\beta_{2}^{-1}\widetilde{p}_{1,\ell_{0}}.$ (A.28)
To compute the limit of $g^{(r)}_{1}(0)$, we first compute it from (A.2) as
$g^{(r)}_{1}(0)$ $=\left(f^{(r)}_{\rho_{1}}(0)+(\rho^{(r)}_{1})^{\ell^{(r)}_{0}}\right)p^{(r)}_{1,0}-\left(f^{(r)}_{1}(0)-\rho^{(r)}_{1}f^{(r)}_{\rho_{1}}(0)\right)\frac{(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}-1}}{1-\rho^{(r)}_{1}}p^{(r)}_{2,\ell^{(r)}_{0}}$
$\quad+\left(f^{(r)}_{1}(0)-f^{(r)}_{\rho_{1}}(0)\right)\frac{\rho^{(r)}_{21}}{(1-\rho^{(r)}_{1})(1-\gamma^{(r)}_{12})}(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}p^{(r)}_{2,\ell^{(r)}_{0}}$
$\quad-\left(f^{(r)}_{\gamma^{-1}_{12}}(0)-f^{(r)}_{\rho_{1}}(0)\right)\frac{(\rho^{(r)}_{1})^{-1}\rho^{(r)}_{21}}{((\rho^{(r)}_{1}\gamma^{(r)}_{12})^{-1}-1)(1-\gamma^{(r)}_{12})}(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}p^{(r)}_{2,\ell^{(r)}_{0}}.$ (A.29)
To proceed to further compute, we note that
$\lim_{r\downarrow 0}rf^{(r)}_{1}(0)=\lim_{r\downarrow 0}r\ell^{(r)}_{0}=\ell_{0},$
$\lim_{r\downarrow 0}rf^{(r)}_{\rho_{1}}(0)=\lim_{r\downarrow 0}r\frac{1-(\rho^{(r)}_{1})^{\ell^{(r)}_{0}}}{1-\rho^{(r)}_{1}}=\beta_{1}^{-1}(1-e^{-\beta_{1}\ell_{0}}),$
$\lim_{r\downarrow 0}f^{(r)}_{\gamma_{12}^{-1}}(0)(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}=\lim_{r\downarrow 0}\frac{1-(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}}{1-\gamma^{(r)}_{12}}\gamma^{(r)}_{12}=\frac{c_{1}\mu}{\lambda_{2}}=\rho_{21}^{-1}.$
Then, we have, using Lemma A.1,
$\lim_{r\downarrow 0}\left(f^{(r)}_{\rho_{1}}(0)+(\rho^{(r)}_{1})^{\ell^{(r)}_{0}}\right)p^{(r)}_{1,0}=\beta_{1}^{-1}(1-e^{-\beta_{1}\ell_{0}})\lim_{r\downarrow 0}\frac{1}{r}p^{(r)}_{1,0}$
[…]